Determine the mass of 20.0 cubic feet of aluminum (density=2.70g/cm^3)
1
Expert's answer
2018-03-20T08:30:16-0400
1 cubic foot = 28.3168466 liters

20 cubic feet = 20*28.3168466 = 566.336932 liters = 566336.932 cm3

m = 2.7g/cm3*566336.932cm3= 1529109.72 g = 1529.11 kg